OpenStudy (anonymous):

factor the algebraic expression below in terms of a single trigonometric function help? sin^2x+sinx-2

OpenStudy (jdoe0001):

sin(x) -2, or sin(x-2)?

OpenStudy (anonymous):

no its (sin^2)x + sin(x) -2 kinda if you look at it like that

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ok.... take a peek at the expression \(\bf sin^2(x)+sin(x)-2\\ \text{let's make sin(x) = a}\\ a^2+a-2\\ \text{now you have a quadratic function}\)

OpenStudy (anonymous):

so now If I solve for "a" that will be my answer? @jdoe0001

OpenStudy (jdoe0001):

so you need 2 numbers, factors of "2" added together should give 1, for 1a middle term multiplied together should give -2, for -2 the constant yes

OpenStudy (anonymous):

so would my answer be....sin^3-2? @jdoe0001

OpenStudy (jdoe0001):

ahemm... no :/ on a quadratic usually you'd factor it to 2 binomials 2 factors of 2 that add to 1 and multiply to -2 let's see +2 * -1 = -2 +2 +(-1) = 1 low and behol! so your binomials will be (a + 2)(a - 1) now if we change the "a" back to sin(x) \(\bf ( sin(x) +2 ) ( sin(x) - 1)\)

OpenStudy (anonymous):

oh ok great haha, @jdoe0001 , so then multiply and I have my answer?

OpenStudy (jdoe0001):

well, no multiplication for this instance, all they want is \(\bf \text{ in terms of a single trigonometric function }\)

OpenStudy (anonymous):

that's just in form of sin(x)-2 ? @jdoe0001

OpenStudy (anonymous):

oh no then I would just leave it in the form you gave me ?

OpenStudy (jdoe0001):

yes, just leave in the forms of the 2 binomial factors, since is all that's being required :)

OpenStudy (anonymous):

super, thank you so much sorry I was trying to refigure this out, but anyhow thanks a billon!!! @jdoe0001

OpenStudy (jdoe0001):

yw